# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ I-F Contributions of the Article
Figure 6: 6G vision: smart wireless world via 6G-enabled wireless intelligence.
II 6G Vision
@@
In this section, our 6G vision is introduced. First, the big picture of 6G is discussed. Then, the focus is shifted to the main details. Specifically, a brief introduction is provided to the fundamental elements, disruptive applications, key use cases, target performance requirements, potential technologies, and defining features. These topics are discussed in detail in their dedicated sections later in the paper. It is worth mentioning that our 6G vision is well aligned with ITU-R’s one, although more far-reaching with a broader perspective.
@@
  • Big Picture: Based on the current trends, we envision that the next major disruption in mobile communications will be the 6G-enabled mobile intelligence. Mobile intelligence refers to the fusion of wireless, AI, and IoE technologies at all levels of society. Mobile intelligence has the potential to make anything connected, smart, and aware of the surrounding environment. Consequently, mobile intelligence will revolutionize the design, operation, interactions, and use of devices, systems, and applications. We envisage that 6G can ultimately become the Intelligent Network of Everything and serve as an enabling platform for the mobile intelligence. The 6G-enabled mobile intelligence will create a smart wireless world, where there are unprecedented opportunities to produce greater value for the benefit of people, society, and the world in general. Eventually, mobile intelligence will penetrate all walks of life and become an essential part of the future society.
@@
  • Fundamental Elements: To provide a major disruption, 6G needs to be based on three profound elements: wireless, AI, and IoE. Wireless consists of connectivity, computation, sensing, and energy dimensions. AI refers to the extensive use of AI at all levels of the 6G ecosystem, including the network core, network edges, air interface, devices, services, and applications. IoE refers to the massive number of network-connected objects, such as sensors, devices, machines, vehicles, drones, robots, etc. Further details on the fundamental elements are provided in Section III.
@@
    For smart environments, 6G needs to provide customized wireless network services with IoE communications, beyond-communication capabilities, and secure network solutions. Connected autonomous systems refer to entities that can operate independently without human involvement, relying on versatile technologies, such as AI, control, sensing, positioning, and connectivity. We define three systems: connected autonomous vehicle systems, connected autonomous aerial vehicle systems, and connected autonomous robotic systems. Such systems set stringent performance requirements for 6G, particularly in terms of reliability, availability, latency, and mobility. A detailed discussion of 6G applications is provided in Section IV. The 6G vision of the smart wireless world with the aforementioned applications is illustrated in Figure 6.
  • Key Use Cases: Since 6G is expected to significantly expand the capabilities of mobile networks, we introduce five communication-oriented and three beyond-communication-oriented use cases accordingly. In the communication domain, there are six main performance dimensions that must be optimized for 6G, i.e., capacity, latency, reliability, density, coverage, and mobility. The corresponding use cases are ultra-broadband multimedia communications, extreme time-sensitive and mission-critical communications, ultra-massive communications, global-scale communications, and hyper-mobility communications. In addition to the enhanced communication features, 6G aims to extend its capabilities beyond communication. The corresponding use cases include network intelligence, network sensing, and network energy. Further details on these use cases are provided in Section V.
  • Performance Requirements: To support diverse use cases and applications, the performance requirements of 6G need to be pushed to their limits. ITU-R has introduced example target values for nine main performance metrics: peak rate (50/100/200 Gbit/s), user experienced rate (300/500 Mbit/s), spectral efficiency (1.5X/3X), area traffic capacity (30/50 Mbit/s/m2), latency (0.1–1 ms), reliability (1-10-${}^{\text{-}5}$–1-10-${}^{\text{-}7}$), connection density (106–108/km2), mobility (500–1000 km/h), and positioning accuracy (1–10 cm). Compared to the targets proposed in the 6G literature, ITU-R’s values are rather moderate. A more detailed discussion on the target requirements is provided in Section VI.
@@
    In the network intelligence domain, AI/ML plays a revolutionary role by empowering the network core, edges, and air interface. Integrating communication, computation, sensing, and energy greatly expands the capabilities of mobile networks. Green communication and networking provide significant ecological benefits, making 6G more energy-efficient and sustainable. Advanced device-to-device (D2D), V2X, and cellular-connected UAV communications extend the support for the IoE-type connectivity in the 6G era. The main service-oriented technologies are private networks, which significantly broaden the service and business opportunities of mobile networks in diverse vertical domains. A holistic network security architecture is vital for the success of 6G, providing trustworthy, secure, and privacy-protected use. In Section VIII, the potential 6G technologies are reviewed in a tutorial manner.
  • Defining Features: We introduce 12 features that define the essence of 6G: extreme capacity &\&& performance, ultra-flexible &\&& agile, highly intelligent &\&& aware, ubiquitously available &\&& reliable, truly green &\&& sustainable, and thoroughly secure &\&& trustworthy. The core of 6G will be extreme performance, especially in terms of capacity, latency, reliability, density, coverage, and mobility. 6G needs to be flexible and agile to adapt to versatile wireless environments and application scenarios. A high level of intelligence and awareness requires pervasive AI and accurate sensing, providing many benefits from efficient network optimization and increased automation to object recognition and velocity estimation. To support diverse services, 6G must be broadly available and reliable. To obtain significant ecological benefits, sustainability and greenness need to be integral parts of 6G. Since 6G will be a pivotal part of the future society, it needs to be secure and trustworthy. Further discussion on the defining features can be found in Section IX.
@@ III Fundamental Elements of 6G
  • Wireless: In the context of 6G, wireless refers to a variety of technologies including connectivity, sensing, positioning, and energy. Extreme connectivity lays the foundation for 6G and its numerous applications. There are six main performance dimensions that need to be pushed to the extreme levels, i.e., capacity, latency, reliability, density, coverage, and mobility. Consequently, 6G is expected to support a vast range of communication scenarios, including ultra-fast mobile internet, time-sensitive and mission-critical connectivity, ultra-dense IoT, global coverage, and hyper-mobility communications. Beyond-communication wireless technologies, such as high-resolution sensing, accurate positioning, and wireless energy, expand the capabilities of mobile networks beyond the current horizon. For example, mobile networks can provide novel services, like object recognition/identification, location-based services, and wireless powering of lightweight IoT devices.
@@
IV Disruptive Applications for 6G
@@
Figure 8: Disruptive applications for 6G.
IV-A Human-Machine Interactions
@@ IV-A1 Metaverse
The term ”metaverse” was first used in science fiction in 1992 [98]. In the 2010s, the term metaverse re-emerged and evolved toward the current vision. Simultaneously, metaverse began to gain more attention in science, becoming a popular topic on the verge of the 2020s. The COVID-19 pandemic has further increased interest in the metaverse since many physical activities have moved to the virtual space [98]. The metaverse is widely recognized in the technology industry. Major technology companies are investing in the development of practical metaverse applications. For example, Facebook was renamed Meta in 2021 to better describe the company’s vision [98]. In practice, there exist some applications that are considered rudimentary versions of the metaverse, mainly in gaming. Second Life, Roblox, and Fortnite are examples of such gaming platforms [97]. In the standardization domain, there exist two metaverse-related standards [97], i.e., ISO/IEC 23005 and IEEE 2888. ISO/IEC 23005 focuses on standardizing the interfaces between the real-virtual and virtual-virtual worlds. IEEE 2888 complements ISO/IEC 23005 by standardizing the synchronization between the real and virtual worlds. Further details on the fundamentals, latest advances, and future challenges of the metaverse can be found in recent survey papers: generic [99, 105, 97, 98, 106, 107] and 6G-related [100, 101, 102, 103, 104].
IV-A2 Extended Reality
@@
IV-A3 Holographic-type Communication
Holographic-type communication refers to the transfer of holograms from a source to a (remote) destination via wireless/wired networks [113]. A hologram is a realistic 3D representation of an object based on the depth and parallax information [113]. Unlike a traditional 3D image, a hologram changes according to the viewer’s position [114]. Holography refers to a technology creating holograms [113]. Specifically, a camera array is placed around a 3D object to collect image information from multiple angles and views. This information is then processed to generate a 3D hologram. To summarize an end-to-end holographic system, holograms are generated using holography (or stored) at the source, transferred via holographic-type communications to the destination, and finally presented by a holographic display [113]. There are three main types of displays: XR HMDs, volumetric displays, and light-field displays [113]. XR HMDs are close to the user’s eyes, with two view angles and looser requirements for data transfer. Volumetric displays are suitable for mobile devices due to their small size and limited viewing angles. Light-field displays have the highest quality with thousands of view angles, multi-user support, and extremely high requirements for holographic-type communications.
@@ IV-A5 Tactile Internet
The concept of Tactile Internet was first introduced in 2014 [130]. In 2016, a working group was established for the first standards (i.e., IEEE P1918.1) related to Tactile Internet [128]. The main objective of IEEE P1918.1 was to specify the architectural basis for Tactile Internet. In the research domain, 5G was first seen as an enabler for Tactile Internet due to its support for URLLC with a latency of up to 1 ms [128]. However, 5G was not designed to support Tactile Internet and its demanding applications. Hence, 5G is not able to serve as an enabling platform for Tactile Internet in practice. Despite the lack of broad support, 5G still has the potential to enable light versions of tactile applications and pave the way for 6G networks as a true enabler. 6G is expected to provide a fruitful platform for Tactile Internet and enable the broad use of novel real-time interactive applications [131]. 6G’s capabilities go way beyond 5G’s in every aspect, especially in terms of performance, intelligence, flexibility, and security, all of which are essential for realizing Tactile Internet in practice. Further details on Tactile Internet can be found in [128, 129, 131].
IV-B Smart Environments
@@ IV-B1 Smart Society
IV-B2 Smart City
By relying on advanced information and communication technologies, the concept of a smart city aims to improve the city’s management, operation, maintenance, and services to offer diverse advantages and a better quality of life for its residents. The synergy among advanced digital technologies provides a fruitful platform to enable smart cities worldwide. Smart cities will be more efficient in numerous ways and will offer a broad range of improved and novel services for their residents. A smart city consists of diverse elements [132, 133], such as smart governance, smart economy, smart environment (nature), smart transportation, smart traffic, smart water, smart waste, smart sanitation, smart schools, smart buildings, smart homes, smart lighting, smart culture, smart sports, smart activities, smart parking, smart shopping, and so on. The opportunities, challenges, priorities, and realizations of smart cities depend on many different factors, such as the population, location, climate, environment, and area of the city [132, 133]. Thus, each city is different, and smart elements need to be designed accordingly. Currently, numerous smart city plans and projects are ongoing around the world [132, 133]. It is envisioned that smart cities will prosper in the 6G era since key technologies have become sufficiently mature to enable such complex entities [133, 134, 135]. Recent information on smart cities is available in [132, 133, 134, 136, 137, 135].
IV-B3 Smart Factory
@@ IV-B4 Smart Home
Smart homes are anticipated to become popular in the 6G era due to the maturation of key technologies, such as intelligent cloud with AI-empowered computing and control, high-accuracy sensing and indoor localization, seamless wireless connectivity and IoT support, as well as data security and privacy. This will open new business opportunities for technology companies (such as mobile network operators, vendors, and service providers) to provide customized and comprehensive smart home solutions with plug-and-play styles. This requires cooperation between technology companies, constructors, and building/homeowners. Further discussions on the smart home concept can be found in [143, 144, 145, 146, 147, 148].
IV-C Connected Autonomous Systems
@@
IV-C1 Connected Autonomous Vehicle Systems
@@
IV-C2 Connected Autonomous Aerial Vehicle Systems
@@
IV-C3 Connected Autonomous Robotic Systems
@@
V Key Use Cases for 6G
6G is expected to greatly expand the capabilities of mobile networks by enhancing ”traditional” communication features and introducing new beyond-communication ones. Following this evolution, we introduce five communication-oriented and three beyond-communication-oriented use cases, supporting a wide range of new 6G-level application scenarios. These use cases are summarized in Figure 9.
V-A Communication-Oriented Use Cases

### v3

@@ II-C DISCUSSION
Figure 6: 6G vision: smart wireless world via 6G-enabled mobile intelligence.
III 6G VISION
@@
In this section, our 6G vision is introduced. First, the big picture of 6G is discussed. Then, the focus is shifted to the main details. Specifically, a brief introduction is provided to the fundamental elements, disruptive applications, key use cases, target performance requirements, potential technologies, and defining features. These topics are discussed in detail in their dedicated sections later in the paper. It is worth mentioning that our 6G vision is well aligned with that of ITU-R, although more far-reaching with a broader perspective.
@@
  • Big Picture: Based on the current trends, we envision that the next major disruption in mobile communications will be the 6G-enabled mobile intelligence. Mobile intelligence refers to the fusion of wireless, AI, and IoE technologies at all levels of the society. Mobile intelligence has the potential to make anything connected, smart, and aware of the surrounding environment. Consequently, mobile intelligence will revolutionize the design, operation, interactions, and use of devices, systems, and applications. We envisage that 6G will evolve toward the Intelligent Network of Everything, while serving as a fruitful platform for mobile intelligence to grow toward its potential. The 6G-enabled mobile intelligence will create a smart wireless world, where there are unprecedented opportunities to produce greater value for the benefit of people, society, and the world in general. Eventually, mobile intelligence will penetrate all walks of life and become an essential part of the future society.
@@
  • Fundamental Elements: To provide a major disruption, 6G needs to be based on three profound elements: wireless, AI, and IoE. Wireless consists of communication, sensing, and energy dimensions. AI refers to the extensive use of AI at all levels of the 6G ecosystem, including the network core, network edges, air interface, devices, services, and applications. IoE refers to the massive number of network-connected objects, such as sensors, devices, machines, vehicles, drones, robots, etc. Further details on the fundamental elements are provided in Section IV.
@@
    For smart environments, 6G needs to provide customized wireless network services with IoE communications, beyond-communication capabilities, and secure network solutions. Connected autonomous systems refer to entities that can operate independently without human involvement, relying on versatile technologies, such as AI, control, sensing, positioning, and connectivity. We define three systems: connected autonomous vehicle systems, connected autonomous aerial vehicle systems, and connected autonomous robotic systems. Such systems set stringent performance requirements for 6G, particularly in terms of reliability, availability, latency, and mobility. A detailed discussion of 6G applications is provided in Section V. The 6G vision of the smart wireless world with the aforementioned applications is illustrated in Figure 6.
  • Key Use Cases: Since 6G is expected to significantly expand the capabilities of mobile networks, we introduce five communication-oriented and three beyond-communication-oriented use cases accordingly. In the communication domain, there are six main performance dimensions that must be optimized for 6G, i.e., capacity, latency, reliability, density, coverage, and mobility. These performance dimensions are among the list of IMT-2030 capabilities introduced by ITU-R in [20]. The corresponding use cases are ultra-broadband multimedia communications, extreme time-sensitive and mission-critical communications, ultra-massive communications, global-scale communications, and hyper-mobility communications. In addition to the enhanced communication features, 6G aims to extend its capabilities beyond communication. The corresponding use cases include network intelligence, network sensing, and network energy. Further details on these use cases are provided in Section VI.
  • Performance Requirements: To support diverse use cases and applications, the performance requirements of 6G need to be pushed to their limits. ITU-R has introduced example target values for nine main performance metrics: peak rate (50/100/200 Gbit/s), user experienced rate (300/500 Mbit/s), spectral efficiency (1.5X/3X), area traffic capacity (30/50 Mbit/s/m2), latency (0.1–1 ms), reliability (1-10-${}^{\text{-}5}$–1-10-${}^{\text{-}7}$), connection density (106–108/km2), mobility (500–1000 km/h), and positioning accuracy (1–10 cm). Compared to the targets proposed in the 6G literature, the values of ITU-R are rather moderate. A more detailed discussion on the target requirements is provided in Section VII.
@@
    In the network intelligence domain, AI/ML plays a revolutionary role by empowering the network core, edges, and air interface. Integrating communication, computation, sensing, and energy greatly expands the capabilities of mobile networks. Green communication and networking provide significant ecological benefits, making 6G more energy-efficient and sustainable. Advanced device-to-device (D2D), V2X, and cellular-connected UAV communications extend the support for the IoE-type connectivity in the 6G era. The main service-oriented technologies are private networks, which significantly broaden the service and business opportunities of mobile networks in diverse vertical domains. A holistic network security architecture is vital for the success of 6G, providing trustworthy, secure, and privacy-protected use. In Section IX, the potential 6G technologies are reviewed in a tutorial manner.
  • Defining Features: We introduce 12 features that define the essence of 6G: extreme capacity &\&& performance, ultra-flexible &\&& agile, highly intelligent &\&& aware, ubiquitously available &\&& reliable, truly green &\&& sustainable, and thoroughly secure &\&& trustworthy. The core of 6G will be extreme performance, especially in terms of capacity, latency, reliability, density, coverage, and mobility. 6G needs to be flexible and agile to adapt to versatile wireless environments and application scenarios. A high level of intelligence and awareness requires pervasive AI and accurate sensing, providing many benefits from efficient network optimization and increased automation to object recognition and velocity estimation. To support diverse services, 6G must be broadly available and reliable. To obtain significant ecological benefits, sustainability and greenness need to be integral parts of 6G. Since 6G will be a pivotal part of the future society, it needs to be secure and trustworthy. Further discussion on the defining features can be found in Section X.
@@ IV FUNDAMENTAL ELEMENTS OF 6G
  • Wireless: In the context of 6G, wireless refers to a variety of technologies including communication, sensing, positioning, and energy. Extreme communication lays the foundation for 6G and its numerous applications. In the communication domain, there are six main performance dimensions that need to be pushed to the extreme levels, i.e., capacity, latency, reliability, density, coverage, and mobility. Consequently, 6G is expected to support a vast range of communication scenarios, including ultra-fast mobile internet, time-sensitive and mission-critical connectivity, ultra-dense IoT, global coverage, and hyper-mobility communications. Beyond-communication wireless technologies, such as high-resolution sensing, accurate positioning, and wireless energy, expand the capabilities of mobile networks beyond the current horizon. For example, mobile networks can provide novel services, like object recognition/identification, location-based services, and wireless powering of lightweight IoT devices.
@@
Figure 8: Disruptive applications for 6G.
V DISRUPTIVE APPLICATIONS FOR 6G
@@
V-A HUMAN-MACHINE INTERACTIONS
@@ V-A1 METAVERSE
The term ”metaverse” was first used in science fiction in 1992 [117]. In the 2010s, the term metaverse re-emerged and evolved toward the current vision. Simultaneously, metaverse began to gain more attention in science, becoming a popular topic on the verge of the 2020s. The COVID-19 pandemic has further increased interest in the metaverse since many physical activities have moved to the virtual space [117]. The metaverse is widely recognized in the technology industry. Major technology companies are investing in the development of practical metaverse applications. For example, Facebook was renamed Meta in 2021 to better describe the vision of the company [117]. In practice, there exist some applications that are considered rudimentary versions of the metaverse, mainly in gaming. Second Life, Roblox, and Fortnite are examples of such gaming platforms [116]. In the standardization domain, there exist two metaverse-related standards [116], i.e., ISO/IEC 23005 and IEEE 2888. ISO/IEC 23005 focuses on standardizing the interfaces between the real-virtual and virtual-virtual worlds. IEEE 2888 complements ISO/IEC 23005 by standardizing the synchronization between the real and virtual worlds. Further details on the fundamentals, latest advances, and future challenges of the metaverse can be found in recent survey papers: generic [118, 124, 116, 117, 125, 126] and 6G-related [119, 120, 121, 122, 123].
V-A2 EXTENDED REALITY
@@
V-A3 HOLOGRAPHIC-TYPE COMMUNICATION
Holographic-type communication refers to the transfer of holograms from a source to a (remote) destination via wireless/wired networks [132]. A hologram is a realistic 3D representation of an object based on the depth and parallax information [132]. Unlike a traditional 3D image, a hologram changes according to the position of the viewer [133]. Holography refers to a technology creating holograms [132]. Specifically, a camera array is placed around a 3D object to collect image information from multiple angles and views. This information is then processed to generate a 3D hologram. To summarize an end-to-end holographic system, holograms are generated using holography (or stored) at the source, transferred via holographic-type communications to the destination, and finally presented by a holographic display [132]. There are three main types of displays: XR HMDs, volumetric displays, and light-field displays [132]. XR HMDs are close to the eyes of the user, with two view angles and looser requirements for data transfer. Volumetric displays are suitable for mobile devices due to their small size and limited viewing angles. Light-field displays have the highest quality with thousands of view angles, multi-user support, and extremely high requirements for holographic-type communications.
@@ V-A5 TACTILE INTERNET
The concept of Tactile Internet was first introduced in 2014 [149]. In 2016, a working group was established for the first standards (i.e., IEEE P1918.1) related to Tactile Internet [147]. The main objective of IEEE P1918.1 was to specify the architectural basis for Tactile Internet. In the research domain, 5G was first seen as an enabler for Tactile Internet due to its support for URLLC with a latency of up to 1 ms [147]. However, 5G was not designed to support Tactile Internet and its demanding applications. Hence, 5G is not able to serve as an enabling platform for Tactile Internet in practice. Despite the lack of broad support, 5G still has the potential to enable light versions of tactile applications and pave the way for 6G networks as a true enabler. 6G is expected to provide a fruitful platform for Tactile Internet and enable the broad use of novel real-time interactive applications [150]. The capabilities of 6G go way beyond to those of 5G in every aspect, especially in terms of performance, intelligence, flexibility, and security, all of which are essential for realizing Tactile Internet in practice. Further details on Tactile Internet can be found in [147, 148, 150].
V-B SMART ENVIRONMENTS
@@ V-B1 SMART SOCIETY
V-B2 SMART CITY
By relying on advanced information and communication technologies, the concept of a smart city aims to improve the management, operation, maintenance, and services to offer diverse advantages and a better quality of life for its residents. The synergy among advanced digital technologies provides a fruitful platform to enable smart cities worldwide. Smart cities will be more efficient in numerous ways and will offer a broad range of improved and novel services for their residents. A smart city consists of diverse elements [151, 152], such as smart governance, smart economy, smart environment (nature), smart transportation, smart traffic, smart water, smart waste, smart sanitation, smart schools, smart buildings, smart homes, smart lighting, smart culture, smart sports, smart activities, smart parking, smart shopping, and so on. The opportunities, challenges, priorities, and realizations of smart cities depend on many different factors, such as the population, location, climate, environment, and area of the city [151, 152]. Thus, each city is different, and smart elements need to be designed accordingly. Currently, numerous smart city plans and projects are ongoing around the world [151, 152]. It is envisioned that smart cities will prosper in the 6G era since key technologies have become sufficiently mature to enable such complex entities [152, 153, 154]. Recent information on smart cities is available in [151, 152, 153, 155, 156, 154].
V-B3 SMART FACTORY
@@ V-B4 SMART HOME
Smart homes are anticipated to become popular in the 6G era due to the maturation of key technologies, such as intelligent cloud with AI-empowered computing and control, high-accuracy sensing and indoor localization, seamless wireless connectivity and IoT support, as well as data security and privacy. This will open new business opportunities for technology companies (such as mobile network operators, vendors, and service providers) to provide customized and comprehensive smart home solutions with plug-and-play styles. This requires cooperation between technology companies, constructors, and building/homeowners. Further discussions on the smart home concept can be found in [162, 163, 164, 165, 166].
V-C CONNECTED AUTONOMOUS SYSTEMS
@@
V-C1 CONNECTED AUTONOMOUS VEHICLE SYSTEMS
@@
V-C2 CONNECTED AUTONOMOUS AERIAL VEHICLE SYSTEMS
@@
V-C3 CONNECTED AUTONOMOUS ROBOTIC SYSTEMS
@@
TABLE III: Summary of key use cases for 6G
Use Cases Target Environment Requirements Challenges Enablers Applications
Ultra-Broadband Multimedia Communications Immersive multimedia Hotspot to rural Extreme rates &\&& ultra-low latency Extreme requirements THz &\&& UDNs XR &\&& HTC
Extreme Time Sensitive Mission Critical Commun. Applications with tight constraints Industrial &\&& vehicular Extreme reliability &\&& ultra-low latency Conflicting requirements Fast/robust network processes Smart factory &\&& CAVs &\&& tele-surgery
Ultra-Massive Communications Ultra-massive IoT/IoE Industrial &\&& urban &\&& wide-area Ultra-dense connections Network flexibility Flexible network mechanisms Smart factory &\&& smart city
Global-Scale Communications Global coverage Land &\&& air &\&& sea Super-coverage Cost efficiency &\&& 3D network NTNs Global IoT &\&& remote internet
Hyper-Mobility Communications Connectivity at high device velocities Urban to rural Extreme mobility Mobility management Lower spectrum &\&& fast handovers Ultra-high-speed trains
Network Intelligence Intelligence as a service Industrial &\&& organizational High level of intelligence Powerful computing infra Pervasive AI/ML Smart environments
Network Sensing Ubiquitous sensing Industrial &\&& healthcare Sensing capabilities Accurate sensing ISAC Detection &\&& tracking
Network Energy Wireless energy services Industrial &\&& private WET capabilities Efficiency &\&& distances Energy beamforming Powering IoT devices
VI KEY USE CASES FOR 6G
6G is expected to greatly expand the capabilities of mobile networks by enhancing ”traditional” communication features and introducing new beyond-communication ones. Following this evolution, we introduce five communication-oriented and three beyond-communication-oriented use cases, supporting a wide range of new 6G-level application scenarios. These use cases are summarized in Figure 9 and Table III.
VI-A COMMUNICATION-ORIENTED USE CASES